OpenStudy (anonymous):

Find the roots of the polynomial equation. 2x^3 + 2x^2 – 19x + 20 = 0 3+i/2 , 3-i/2 , -4 -3+2i/2 , -3-2i/2 , 4 -3+i/2 , -3-i/2 , -4 3+2i/2 , 3-2i/2 , 4

OpenStudy (anonymous):

so, you know how to do synthetic division?

OpenStudy (anonymous):

No , im home-schooled . I dont get taught this stuff . its really hard for me because i dont have teacher . . .

OpenStudy (anonymous):

no worries, its about as easy as the rational root test and they go together anyways.

OpenStudy (anonymous):

That's gotta be good , rite ? lol

OpenStudy (anonymous):

to start you need to do the rrt and find all rational roots. In this case there are 12 possible roots, but because its multiple choice we can eliminate the root to be either -4, or 4.

OpenStudy (anonymous):

rite.. :) its a long process...

OpenStudy (anonymous):

now I have to draw

OpenStudy (anonymous):

ok , (:

OpenStudy (anonymous):



OpenStudy (anonymous):

alright so this is synthetic division. You take the polynomial and divide it by the root you want to test. If it is legit, then the remainder (84 above) should be zero. Its not really division though. To start you take the 2 bring is down and times is by 4 giving you 8. Then 8 plus 2 is 10, times by 4 is 40, and so on.

OpenStudy (anonymous):

because the remainder is 84, it is safe to say that 4 is not a root.

OpenStudy (anonymous):

ok, i seee that . so it has to be a or c . .. .

OpenStudy (anonymous):

yep, I tested -4 and it works

OpenStudy (anonymous):

when you do negative 4 you get, _______________________ -4 | 2 2 -19 20 -8 24 -20 ------------------ 2 -6 5 0

OpenStudy (anonymous):

yes i seee. so then how do you find out about the 3+i/2 part ?

OpenStudy (anonymous):

the bottom is important because this is your new polynomial. 2x^2 - 6x + 5=0 Find the roots of that by doing the quadratic equation.

OpenStudy (anonymous):

you know that by heart right? ;)

OpenStudy (anonymous):

dont you do it how you showed me on my last problem ?

OpenStudy (anonymous):

o well, the roots that you have left are not rational.

OpenStudy (anonymous):

so , what does that mean ?

OpenStudy (anonymous):

the rational root test can only test for roots that are whole numbers. Square roots and imaginary numbers cant be done that way.

OpenStudy (anonymous):

So , then how are we supposed to figure the answer out ?

OpenStudy (anonymous):

quadratic equation.

OpenStudy (anonymous):

best tool in finding roots hands down

OpenStudy (anonymous):

dont think , im a slow , lol . but whats that equation. I dont know it by heart .

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{2a }\]

OpenStudy (anonymous):

know how to use it?

OpenStudy (anonymous):

ughhh , i suck at that ^ -____-

OpenStudy (anonymous):

yeah, it generally sucks all around... best thing to do is convert your polynomial into something a bit easier to manage. Like this

OpenStudy (anonymous):

x^2 - 3x + 5/3 = 0 (divide the whole thing by 3)

OpenStudy (anonymous):

3 +/- sqrt(9 - 4(5/3)) all over 2

OpenStudy (anonymous):

do you have to choose plus or minus ?

OpenStudy (anonymous):

nope, they are both included. You will have at least two roots using the quadratic equation.

OpenStudy (anonymous):

i just tried to type that in my calc & it said syntax error.

OpenStudy (anonymous):

whoa, ok. Im back. So I goofed above. The equation is 3 +/- sqrt( 9 - 4(5/2)) all over 2 but if you plug this into your calc it wont work because of the imaginary number. You will have to simplify on paper.

OpenStudy (anonymous):

i just messed up badly . i got 3 +/- (-3-10)

OpenStudy (anonymous):

thats not too bad, you just have to square the 3 and then your all set

OpenStudy (anonymous):

what do you mean ? how ?

OpenStudy (anonymous):

3 +/- (9-10)

OpenStudy (anonymous):

im sooo confused .

OpenStudy (anonymous):

sorry. you will have 3+/- sqrt(9-10) this is 3+/- sqrt(-1) which is 3+/- i

OpenStudy (anonymous):

all of this is divided by 2. SO i think your answer will be c

OpenStudy (anonymous):

i thought it would of been a . im terrible at math . lol

OpenStudy (anonymous):

no worries. Practice is the key.

OpenStudy (anonymous):

well thank you sooo much , i appreciate it (;

OpenStudy (anonymous):

your welcome. Learning math at home without a teach can be hard so hit me up if you need more help.